a body moves with the velocity of 40m/s is brought to rest in 5 seconds . calculate the retardation
1
Expert's answer
2018-04-17T10:09:11-0400
v=v_0-at=0.
The retardation is
a=v_0/t=40/5=8 m/s^2 .